OpenStudy (baby456):

Expontenial function question please help me! Iodine-131 is a radioactive isotope used in the treatment of thyroid conditions.It has a half-life of 8 days/ hsalf life is the amount of daysfor half of the substance to decay. If a patient is iven 20mgof iondine131 how much of the substance will remain in his body after 32 days? a.8mg b.5mg c.2.5mg d.1.25mg

OpenStudy (baby456):

Could you help

OpenStudy (baby456):

@zepdrix

OpenStudy (baby456):

@Hero

OpenStudy (tkhunny):

Please show your work. You were given this problem for a reason. You must have some idea of the required exponential form.

OpenStudy (baby456):

y=20(1-8)^32 is that the right form becayse i think this is expontential decay

OpenStudy (baby456):

I really dont know sometimes my teacher is bad at explaning i dont know how to do it

OpenStudy (tkhunny):

I don't understand that. Where is the number of days? \(y = Ce^{-kt}\) -- t is in days For t = 0, we have C = Initial Population = 20 mg \(y = (20 mg)e^{-kt}\) Half-life of 8 days suggests: \(10 = (20 mg)e^{-k(8)}\) -- The 10 is half of 20. You can now solve for k and complete the rest of the problem.

OpenStudy (baby456):

we were never taught that

OpenStudy (baby456):

i will try

OpenStudy (baby456):

how do you solve it i am stuck

OpenStudy (tkhunny):

Use your logarithms. \(\dfrac{1}{2} = e^{-8k}\) \(ln(1/2) = -8k\) Finish up to find k and you are nearly done.